# FineCops-Ref: A new Dataset and Task for Fine-Grained Compositional Referring Expression Comprehension

### v1

@@ 1 Introduction
Despite significant advancements in multimodal large language models (MLLMs), a critical challenge remains in ensuring these models’ responses are grounded in visual content rather than solely derived from linguistic cues Tong et al. (2024); Zhai et al. (2023); Miyai et al. (2024). Vision-language models often treat language as a bag of words, lacking meaningful engagement with word order, attributes, or relationships Ma et al. (2023); Thrush et al. (2022); Tong et al. (2024); Yuksekgonul et al. (2022), and exhibit poor grounding and spatial reasoning abilities Chen et al. (2024a); Tong et al. (2024); Zhang et al. (2024).
Current evaluation methods utilize Visual Question Answering or Image-Text Retrieval to evaluate the compositional reasoning or grounding abilities of MLLMs. However, these methods provide an indirect assessment of the models’ visual grounding capabilities. In contrast, the Referring Expression Comprehension (REC) task requires the model to directly output the bounding box coordinates based on a given expression, serving as an ideal testing ground for MLLMs.
Recent MLLMs, leveraging substantial grounding data Chen et al. (2023); Wang et al. (2023b, a) and specifically designed visual modules You et al. (2024); Li et al. (2024a), have achieved impressive results on common REC benchmarks like RefCOCO/+/g Yu et al. (2016). However, these benchmarks lack considerations of compositional reasoning, allowing models to perform well without understanding linguistic structure or even without the expression Cirik et al. (2018); Akula et al. (2020). Additionally, current vision-language models struggle with negative samples, where the target object is absent from the image Chen et al. (2020); Kurita et al. (2023); You et al. (2024). This limitation is further exacerbated by the lack of robustness in existing datasets, which fail to provide the necessary complexity and variability to thoroughly evaluate MLLMs.
@@ 2 Related Works
Referring expression comprehension. The REC methods can generally be divided into two categories based on whether or not it uses LLMs: specialist and MLLMs. Specialists typically extract text and image features separately and perform multi-stage fusion Liu et al. (2023c); Yan et al. (2023); Kamath et al. (2021). Their training tasks often include various object location tasks. Recently, Zhao et al. (2024a) achieved excellent results on two visual grounding (VG) benchmarks by leveraging hard negative samples in training.
On the other hand, MLLMs directly input the projected visual features into the LLM. Recent methods aim to enhance grounding capabilities in MLLMs through dataset construction with coordinate information and additional visual modules. Common methods for datasets include transforming traditional visual datasets into an instruction-following format using templates Li et al. (2024b); Pramanick et al. (2023); Wang et al. (2023b), correlating object coordinates with existing captions Peng et al. (2024); Qi et al. (2024), and using GPT to generate question-and-answer pairs based on images, object coordinates, and captions You et al. (2024). The All-Seeing Project Wang et al. (2024) has recently introduced a new dataset (AS-1B) using a scalable data engine that incorporates human feedback and efficient models in the loop.
In terms of visual modules, some methods integrate additional visual components, such as GLaMM Rasheed et al. (2024) and LLaVA-Grounding Zhang et al. (2023), while others extract regional features to use as additional inputs Ma et al. (2024); Shao et al. (2024); You et al. (2024); Li et al. (2024a).
@@
For REC tasks, Akula et al. (2020) critically examined RefCOCOg, showing that 83.7% of test instances do not require reasoning on linguistic structure, and proposed the Ref-Adv dataset, which perturbs original expressions to refer to different target objects.
CLEVR-Ref+ Liu et al. (2019) is a synthetic dataset emphasizing relationships, attributes, and linguistic logic. Cops-Ref Chen et al. (2020) and Ref-Reasoning Yang et al. (2020) use GQA scene graphs Hudson and Manning (2019) and rule-based methods to create large-scale compositional referring expression comprehension datasets in real-world scenarios. Cops-Ref additionally added distracting images based on attributes, relationships, and target names. GITM-MR Wu et al. (2023) explores mismatched relationship in the REC task. RefEgo Kurita et al. (2023) and OmniLabel Schulter et al. (2023) consider out-of-distribution scenarios where referred targets do not exist in the image.
@@ 3 FineCops-Ref
FineCops-Ref includes both positive and negative data. Figure 1 illustrates the data construction pipeline.
@@ 3.1 Creating Positive Data
Path Generation. We employ image scene graphs from GQA Hudson and Manning (2019) for path generation. The scene graphs contain detailed information about objects, attributes, and relations. To ensure accuracy, we first filter the objects based on their suitability as target or related objects. We leverage annotations from InstInpaint Yildirim et al. (2023) and applying additional filters such as keywords and object size.
Next, we generate several paths for each of the filtered objects, as show in Figure 1(a). To eliminate any ambiguity, we utilize unique attributes or relations to identify the target object that share the same category as other objects in the image. and we make sure that every generated path are unique.
Data categorization. We categorize positive expressions into three difficulty levels, depending on the complexity of fine-grained reasoning. Level 1 indicates that there are no objects in the image with the same category as the target object. In this case, model can locate the target without requiring contextual understanding. Level 2 signifies the presence of an object with the same name as the target in the image, and the target can be distinguished through one unique attribute or relation. Level 3 require at least two or more relationships and attribute information. The difficulty level is established based on the intricacy of fine-grained reasoning, rather than the complexity of the textual description.
@@
Cops-ref 12586 - 37758
Ref-adv 9602 3704 -
Ours 9605 9814 8507
Table 1: Comparison between the proposed benchmark and other REC benchmarks. Unconstrained indicates the final expression is not constrained by the templates. Cops. indicates fine-grained compositional reasoning. On the right hand side, the test set count of each benchmark is listed.
Expression Generation. We first employ a data balancing technique that takes into account the proportion of each type, effectively minimizing bias in the scene graph. Subsequently, the generated paths are substituted into predefined templates to generate reference expressions. We detail the predefined templates in Appendix A.1.
To further augment the naturalness and diversity of these expressions, we leverage GPT-3.5-turbo to rewrite the referring expressions. By incorporating well-designed instruction and examples, we are able to achieve a more expansive range of linguistically varied and authentic expressions. Prompts used to rewrite are listd in Appendix A.4.
Human Filter. Owing to the inherent constraints of the scene graph annotation information, the data pertaining to levels 2 and 3 may contain inaccuracies, leading to non-uniqueness in the targets referenced. To address this, human annotators filtered this portion of the data manually. Details refer to Appendix A.5.
@@ 3.2 Generating Negative Data
To conduct a thorough and systematic assessment of the REC in existing MLLMs, we generate hard negatives from both textual and visual sources. Like positive data, negative data are categorized into different levels based on the difficulty. Level 1 signifies alterations made to the target object in the negative data, which are relatively straightforward for the model to identify. Level 2 involves modifications to the related objects, disrupting the contextual information and posing a greater challenge for existing models to recognize.
@@
Generating Negative Expressions. Our array of negative expressions encompasses a wide range of challenging types. We draw inspiration from CREPE Ma et al. (2023) and SUGARCREPE Hsieh et al. (2023) to consider various forms of hard negatives. In total, FineCops-Ref covers 5 fine-grained types of hard negative expressions. These types can be broadly classified into two categories: replace and swap. Replace involves generating a negative expression by substituting a portion of the original expression, whether it is an object, an attribute, or a relation. During replacement, we tested the output quality of several approaches. Ultimately, we found that LLM Replace performed the best. For more information, please refer to Appendix A.3. We utilize LLM to determine the most appropriate negative word, ensuring that the negative expression is genuinely negative while only slightly deviating from the original expression. On the other hand, swap entails generating a negative expression by interchanging two attributes or objects within the same category. We further employ LLM to rewrite the new expression. Please refer to Appendix A.2 for more details.
@@
Generating Negative Images. We consider the necessity of negative images from the following aspects. First, negative images enables a more thorough assessment of models’ visual parsing capabilities. Additionally, the evaluation conducted by Visualgptscore Lin et al. (2023) suggests that negative expressions may lack plausibility and fluency and can be detected by language prior.
@@
We generate hard negative images that bears slight differences from the original, such as alterations in objects, attributes, or relations. When dealing with simple positional relationships, we employ horizontal flips. For more intricate modifications involving objects and attributes, we utilize PowerPaint Zhuang et al. (2023), an exceptional image inpainting model offering versatility, to perform precise edits on the image. To guide PowerPaint in editing the image, we utilize LLM-generated replacements as textual guides and the bounding box as a mask. Overall, FineCops-Ref encompasses 5 distinct types of challenging negative images. Please refer to Appendix A.2 for more details.
@@
Negative Data Debiasing. During the generation of negative samples, it is inevitable that certain implausible and incoherent expressions, as well as unreasonable and easily distinguishable negative images, may emerge. To ensure the benchmark’s quality, we employed various techniques to filter out these unsuitable samples and further improve the quality of the benchmark. For negative expressions, we employ the Adversarial Refinement technique proposed by SUGARCREPE Hsieh et al. (2023). It helps mitigate biases and unintended artifacts in the dataset.
@@
To ensure the exclusion of inappropriate and excessively unreasonable negative images, we employ a multi-step filtering process. First, we use CLIP Radford et al. (2021) to ensure that the similarity between the negative text and the positive image is lower than the similarity between the positive text and the positive image. Next, we apply a diffusion-generated inspection model, DIRE Wang et al. (2023c), to filter out excessively unnatural images, excluding those with scores exceeding 0.2. Subsequently, we use DINOv2 Oquab et al. (2023) to compute the image-image similarity between the positive and negative images, retaining the one with the highest DINOv2 score from the 10 candidate negative images.
@@
3.3 Statistics
FineCops-Ref consists of 9,605 positive expressions, 9,814 negative expressions, and 8,507 negative images. Table 1 provides a comparison between FineCops-Ref and other visual grounding benchmarks. FineCops-Ref combines the advantages of unconstrained expression, fine-grained compositional reasoning, difficulty level, and hard negatives at both textual and visual levels. Additionally, we partition the training set and validation set simultaneously as in Table 2. For more details, please refer to the Appendix A.2.
@@ 4 Experiment
We evaluates several representative models, including both traditional vision-language models (Specialist) and MLLMs. The models examined in this study include MDETR Kamath et al. (2021), MM-GDINO Zhao et al. (2024b); Liu et al. (2023c), UNINEXT Yan et al. (2023), Shikra Chen et al. (2023), Ferret You et al. (2023), GroundingGPT Li et al. (2024b), Lenna Wei et al. (2023), InternVL Chen et al. (2024b), CogVLM Wang et al. (2023a) and CogCom Qi et al. (2024). We use there open-source checkpoints to evaluate.
We additionaly evaluate the GPT4-VAchiam et al. (2023). Since GPT4-V’s ability to directly output bounding boxes is relatively limited, we use GPT4-V combined with the Set-of-Mark (SoM) Yang et al. (2023) to evaluate its performance. The Model source and implementation details are in Appendix B.
We also test the effectiveness of training with the training dataset constructed using our data generation pipeline. We fine-tuned MM-GDINO-T and CogVLM using the positive data from the constructed training set. In addition, we fine-tuned MM-GDINO-T using the entire training set. The training settings are in Appendix B.
We evaluate the models using Precision@1 for positive data; Recall@1 and AUROC for negative data. Specifically, models like MDETR and Lenna that have dedicated object detection modules can generate multiple detection boxes with associated confidence scores, allowing for direct computation of Recall@1 and AUROC. For models that generate coordinates as the text using an autoregressive approach, we use the probability of the coordinates tokens to calculate confidence Kurita et al. (2023); Mitchell et al. (2023).
@@ 4.2 Evaluation on Positive data
MLLMs exhibit superior reasoning abilities. For levels 2 and 3, models need robust language comprehension due to the presence of many easily confusable objects in the images. However, most models do not demonstrate sufficient capability in this aspect.
Multimodal models based on large language models (LLMs) achieved better results in this regard, demonstrating that MLLMs possess stronger compositional reasoning abilities.
@@ 5.2 Is rewrite useful?
For positive data, using template-generated data always places the subject at the beginning of the sentence and has a very clear linguistic structure, which does not adequately assess the model’s language understanding abilities. For negative data, without rewriting, there are issues with non-fluency and nonsensicalityHsieh et al. (2023), which can be easily detected by text-only models such as Grammar Morris et al. (2020) and Vera Liu et al. (2023b).
@@ A.2 Examples of dataset
Figure 3: Positive expressions of different difficulty levels.
@@
Negative images. Figure 5 illustrates negative images generated by different methods.
Figure 5: Negative images generated by different methods.
Dataset statistics. For positive expressions and negative expressions, we split the dataset into train, test, and val sets. Specifically, positive expressions are classified based on levels, as detailed in Table 9. Negative expressions are classified based on types, as detailed in Table 10. While for negative images, we only generated them in the test set, categorized by type. Refer to Table 11 for more details.

### v2

@@ 1 Introduction
Despite significant advancements in multimodal large language models (MLLMs), a critical challenge remains in ensuring these models’ responses are grounded in visual content rather than solely derived from linguistic cues Tong et al. (2024); Zhai et al. (2023); Miyai et al. (2024). Vision-language models (VLMs) often treat language as a bag of words, lacking meaningful engagement with word order, attributes, or relationships Ma et al. (2023); Thrush et al. (2022); Tong et al. (2024); Yuksekgonul et al. (2022), and exhibit poor grounding and spatial reasoning abilities Chen et al. (2024a); Tong et al. (2024); Zhang et al. (2024).
Current evaluation methods utilize Visual Question Answering or Image-Text Retrieval to evaluate the compositional reasoning or grounding abilities of MLLMs. However, these methods provide an indirect assessment of the models’ visual grounding capabilities. In contrast, the Referring Expression Comprehension (REC) task requires a model to directly output the bounding box coordinates of a target object based on a given language expression, serving as an ideal testing ground for MLLMs.
Recent MLLMs, leveraging substantial grounding data Chen et al. (2023); Wang et al. (2023b, a) and specifically designed visual modules You et al. (2024); Li et al. (2024a), have achieved impressive results on common REC benchmarks like RefCOCO/+/g Yu et al. (2016). However, these benchmarks lack considerations of compositional reasoning, allowing models to perform well without understanding linguistic structure or even without the expression Cirik et al. (2018); Akula et al. (2020). Additionally, current VLMs struggle with negative samples, where the target object is absent from the image Chen et al. (2020); Kurita et al. (2023); You et al. (2024). This limitation is further exacerbated by the lack of robustness in existing datasets, which fail to provide the necessary complexity and variability to thoroughly evaluate MLLMs.
@@ 2 Related Works
Referring expression comprehension. The REC methods can generally be divided into two categories based on whether or not it uses LLMs: specialist and MLLMs. Specialists typically extract text and image features separately and perform multi-stage fusion Liu et al. (2023b); Yan et al. (2023); Kamath et al. (2021). Their training tasks often include various object location tasks. Recently, Zhao et al. (2024a) achieved excellent results on two visual grounding (VG) benchmarks by leveraging hard negative samples in training.
On the other hand, MLLMs directly input the projected visual features into the LLM. Recent methods aim to enhance grounding capabilities in MLLMs through dataset construction with coordinate information and additional visual modules. Common methods for datasets include transforming traditional visual datasets into an instruction-following format using templates Li et al. (2024b); Pramanick et al. (2023); Wang et al. (2023b), correlating object coordinates with existing captions Peng et al. (2024); Qi et al. (2024), and using LLMs to generate grounded question-answer pairs based on images, object coordinates, and captions You et al. (2024); Wang et al. (2024).
In terms of visual modules, some methods integrate additional visual components, such as GLaMM Rasheed et al. (2024) and LLaVA-Grounding Zhang et al. (2023), while others extract regional features as additional inputs Ma et al. (2024); Shao et al. (2024); You et al. (2024); Li et al. (2024a).
@@
For REC tasks, Akula et al. (2020) critically examined RefCOCOg, showing that 83.7% of test instances do not require reasoning on linguistic structure, and proposed the Ref-Adv dataset, which perturbs original expressions to refer to different target objects. CLEVR-Ref+ Liu et al. (2019) is a synthetic dataset emphasizing relationships, attributes, and linguistic logic. Cops-Ref Chen et al. (2020) and Ref-Reasoning Yang et al. (2020) use GQA scene graphs Hudson and Manning (2019) and rule-based methods to create large-scale compositional referring expression comprehension datasets in real-world scenarios. Cops-Ref additionally added distracting images based on attributes, relationships, and target names. GITM-MR Wu et al. (2023) explores mismatched relationship in the REC task. RefEgo Kurita et al. (2023) and OmniLabel Schulter et al. (2023) consider out-of-distribution scenarios where referred targets do not exist in the image.
@@ 3 FineCops-Ref
FineCops-Ref includes both positive and negative data. Figure 1 illustrates the dataset construction pipeline.
@@ 3.1 Creating Positive Data
Path Generation. We employ image scene graphs from GQA Hudson and Manning (2019) for path generation. The scene graphs contain detailed information about objects, attributes, and relations. To ensure accuracy, we first filter the objects based on their suitability as target or related objects. We leverage annotations from InstInpaint Yildirim et al. (2023) and apply additional filters such as keywords and object size. Next, we generate several paths for each of the filtered objects, as show in Figure 1(a). To eliminate ambiguity, we utilize unique attributes or relations to identify the target object that shares the same category as other objects in the image to ensure that every generated path is unique.
Data categorization. We categorize the paths into three difficulty levels based on the complexity of fine-grained reasoning. Level 1 indicates that there are no other objects in the image belonging to the same category as the target object, such as the TV in Figure 1. In this case, the model can locate the target without requiring contextual understanding. Level 2 signifies the presence of another object with the same category as the target in the image, where the target can be distinguished through one unique attribute or relation. Level 3 requires at least two or more relationships and attribute information. The difficulty levels are determined by the intricacy of fine-grained reasoning, rather than the complexity of the textual description.
@@
Cops-ref 12,586 - 37,758
Ref-adv 9,602 3,704 -
Ours 9,605 9,814 8,507
Table 1: Comparison between the proposed benchmark and other REC benchmarks. Unconstrained indicates the final expression is not constrained by the templates. Cops. indicates compositional reasoning. On the right hand side, the test set count of each benchmark is listed.
Expression Generation. To mitigate potential biases in the scene graphs, we first apply frequency-based sampling of relationships, attributes, and object categories along the generated paths. Subsequently, we use predefined templates to generate referring expressions. The details of predefined templates are provided in Appendix A.1.
To further augment the naturalness and diversity of these expressions, we leverage GPT-3.5-turbo to rewrite the referring expressions. By incorporating well-designed instructions and examples, we achieve a wider range of linguistically diverse and natural expressions. The prompts used for rewriting are listed in Appendix A.4.
Human Filter. Due to inherent limitations in the scene graph annotations, the generated paths, particularly for Level 2 and Level 3, may contain inaccuracies, leading to non-unique target references. To address this, human annotators manually filtered the test set. Further details can be found in Appendix A.5.
@@ 3.2 Generating Negative Data
To conduct a thorough and systematic assessment of REC of existing VLMs, we generate hard negatives from both textual and visual sources. Similar to positive data, negative data are categorized into different levels based on difficulty. Negative Level 1 involves alterations to the target object in the expressions, which are relatively straightforward for the model to identify. Level 2 involves modifications to the related objects, disrupting the contextual information and posing a greater challenge for existing models to recognize.
@@
Generating Negative Expressions. Our set of negative expressions encompasses a wide range of challenging types. Inspired by CREPE Ma et al. (2023) and SUGARCREPE Hsieh et al. (2023), we consider various forms of hard negatives. In total, FineCops-Ref covers 5 fine-grained types of hard negative expressions. These types can be broadly classified into two categories: Replace and Swap. The Replace category involves generating negative expressions by substituting a portion of the original expression, whether it is an object, attribute, or relation. We utilize LLM to determine the most appropriate negative word, ensuring that the negative expression is genuinely negative while only slightly deviating from the original expression. We experimented with various replacement methods and found that LLM-based replacements performed the best, as discussed in Appendix A.3. The Swap category entails generating negative expressions by interchanging two attributes or objects within the same category. We further employ LLM to rewrite these new expressions.
@@
Generating Negative Images. We consider the necessity of negative images from the following aspects. First, negative images enables a more thorough assessment of models’ visual parsing capabilities. Additionally, evaluations conducted by Visualgptscore Lin et al. (2023) suggest that negative expressions may lack plausibility and fluency and can be detected by language prior.
@@
We generate hard negative images with subtle differences from the original, such as modifications to objects, attributes, or relations. For simple positional relationships, we employ horizontal flips. For more intricate modifications involving objects and attributes, we utilize PowerPaint Zhuang et al. (2023), an exceptional image inpainting model offering versatility, to perform precise edits on the images. To guide PowerPaint in editing the image, we utilize LLM-generated replacements as textual guides and the bounding boxes as masks. Overall, FineCops-Ref encompasses 5 distinct types of challenging negative images. Further details on the data types, along with example expressions and images, can be found in Appendix A.2.
@@
Negative Data Debiasing. During the generation of negative samples, some implausible or incoherent expressions, as well as unreasonable and easily distinguishable negative images, are inevitable. We employed several techniques to filter out unsuitable samples and improve the quality of the benchmark. For negative expressions, we employ the Adversarial Refinement technique proposed by SUGARCREPE Hsieh et al. (2023), which helps mitigate biases and unintended artifacts in the dataset.
@@
To exclude inappropriate and excessively unreasonable negative images, we employ a multi-step filtering process. First, we use CLIP Radford et al. (2021) to ensure that the similarity between the negative text and the positive image is lower than the similarity between the positive text and the positive image. Next, we apply the diffusion-generated inspection model DIRE Wang et al. (2023c) to filter out excessively unnatural images, excluding those with scores above 0.2. Finally, we use DINOv2 Oquab et al. (2023) to compute the image-image similarity between the positive image and the generated negative images, retaining the candidate negative image with the highest DINOv2 score from a set of 10 candidates, thereby minimizing noise.
@@
3.3 Dataset Statistics
FineCops-Ref consists of 9,605 positive expressions, 9,814 negative expressions, and 8,507 negative images in test set. Table 1 provides a comparison between FineCops-Ref and other visual grounding benchmarks. FineCops-Ref combines the advantages of unconstrained expression, fine-grained compositional reasoning, difficulty level, and hard negatives at both textual and visual levels. Additionally, we partition the training set and validation set simultaneously as in Table 2. For more statistics, please refer to the Appendix A.2.
@@ 4 Experiment
We evaluates several representative models, including both traditional vision-language models (Specialist) and MLLMs. The models examined in this study include MDETR Kamath et al. (2021), MM-GDINO Zhao et al. (2024b); Liu et al. (2023b), UNINEXT Yan et al. (2023), Shikra Chen et al. (2023), Ferret You et al. (2024), GroundingGPT Li et al. (2024b), Lenna Wei et al. (2023), InternVL Chen et al. (2024b), CogVLM Wang et al. (2023a) and CogCom Qi et al. (2024). We use there open-source checkpoints to evaluate. We additionaly evaluate the GPT4-VAchiam et al. (2023). Since GPT4-V’s ability to directly output bounding boxes is relatively limited, we use GPT4-V combined with the Set-of-Mark (SoM) Yang et al. (2023) to evaluate its performance. The Model source and implementation details are in Appendix B.
We also test the effectiveness of training with the training dataset constructed using our data generation pipeline. We fine-tuned MM-GDINO-T and CogVLM using the positive data from the constructed training set. In addition, we fine-tuned MM-GDINO-T with the entire training set. The training settings are detailed in Appendix B.
We evaluate the models using Precision@1 for positive data, and Recall@1 and AUROC for negative data; the AUROC results can be found in Appendix C. Specifically, models like MDETR and Lenna that have dedicated object detection modules can generate multiple detection boxes with associated confidence scores, allowing for direct computation of Recall@1 and AUROC. For models that generate coordinates as text using an autoregressive approach, we use the probability of the coordinate tokens to calculate confidence Kurita et al. (2023); Mitchell et al. (2023).
@@ 4.2 Evaluation on Positive data
MLLMs exhibit superior reasoning abilities. For levels 2 and 3, models need robust language comprehension due to the presence of many easily confusable objects in the images. However, most models do not demonstrate sufficient capability in this aspect. Multimodal models based on large language models (LLMs) achieved better results in this regard, demonstrating that MLLMs possess stronger compositional reasoning abilities.
@@ 5.2 Is rewrite useful?
For positive data, using template-generated data always places the subject at the beginning of the sentence and has a very clear linguistic structure, which does not adequately assess the model’s language understanding abilities. For negative data, without rewriting, there are issues with non-fluency and nonsensicality Hsieh et al. (2023), which can be easily detected by text-only models such as Grammar Morris et al. (2020).
@@ A.1 Predefined templates
Figure 3: Positive expressions of different difficulty levels.
Figure 5: Negative images generated by different methods.
@@ A.2 Examples of dataset
Negative images. Figure 5 illustrates negative images generated by different methods. Dataset statistics. For positive expressions and negative expressions, we split the dataset into train, test, and val sets. Specifically, positive expressions are classified based on levels, as detailed in Table 9. Negative expressions are classified based on types, as detailed in Table 10. While for negative images, we only generated them in the test set, categorized by type. Refer to Table 11 for more details.